Find all solutions for:
4 sin^2 x - 3 = 0

Answer by lwsshak3(11628)

Find all solutions for:
4 sin^2 x - 3 = 0
sin^2x=3/4
take sqrt of both sides
sinx=√3/2
x=π/3+2πk, 2π/3+2πk, k=integer